OpenStudy (barrelracing):

suppose y varies directly with x. write a direct variation equation the relates x and y. then find the value of y when x =16 y=5 when x=2

rishavraj (rishavraj):

For direct variation y = kx k is the propotionality constany

OpenStudy (barrelracing):

@bry856

jimthompson5910 (jim_thompson5910):

y = k*x y = k*2 ... replace x with 2 5 = k*2 ... replace y with 5 solve for k to get k = _________ (fill in the blank)

rishavraj (rishavraj):

After getting k ...plug the value of k and x in the equation and get y

OpenStudy (barrelracing):

y=40

jimthompson5910 (jim_thompson5910):

you should have found that k = 2.5 y = 2.5x when x = 16, y = 2.5*16 = 40 so you have the correct y value

OpenStudy (barrelracing):

how do I find the direct variation that related x and y?

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (barrelracing):

it says write a direct variation equation that relates x and y. this is still the same problem.

jimthompson5910 (jim_thompson5910):

so it just wants the equation with x and y in it

jimthompson5910 (jim_thompson5910):

the k will be replaced with the numeric value you found

OpenStudy (barrelracing):

so y=40x?

jimthompson5910 (jim_thompson5910):

So you can either say \[\Large y = \frac{5}{2}x\] or \[\Large y = 2.5x\]

jimthompson5910 (jim_thompson5910):

I started with y = kx then replaced k with 5/2 or 2.5

OpenStudy (barrelracing):

ohh ok I get it now thank you

jimthompson5910 (jim_thompson5910):

np